# The Case for ‘Thick Evaluations’ of Cultural Representation in AI

### v1

@@ 5. Discussion
In this discussion section, we reflect on three implications of our findings for the enterprise of AI evaluations: (1) the need to reconsider the construct of representation underpinning AI evaluations and creating congruence between concept and measurement, (2) the importance of co-constructing methods for evaluating representation with communities, and (3) the tensions inherent in attempts to “thicken” evaluation practices for cultural representation in AI.
5.1. Thick Methods for Thick Constructs of Cultural Representation
@@
Our findings highlight the urgent need for a critical re-evaluation of how we conceptualize and operationalize representation in AI evaluations. Instead of representation being a singular fixed concept or objective truth, participants unpacked it in situated and negotiated ways into multiple granular categories. Yet, as Chasalow and Levy (2021) argue, the conceptualization of representation is often taken for granted in AI evaluations, as it is a “suitcase word” encompassing a multitude of meanings and lineages, including political, statistical, and here, cultural. Our participants empirically demonstrated that that what gets bundled into this suitcase word are multiple unobservable constructs for cultural representation, each requiring different approaches to measurement and metrics. Our findings thus suggest that AI evaluations typically conflate two constructs of representation in one—one that is thinner, which focuses on empirical accuracy of physical worlds, and one that is thicker, such as “coherence,” which reflects the internal social consistency and plausibility of a depicted scene.
@@
While thin evaluations may be suitable for assessing the correspondence of “thinner” constructs like factual incorrectness with a base empirical reality or ground truth (e.g., landmarks), thicker constructs, such as those related to cultural norms, social relations, and power dynamics, necessitate thick evaluations that can capture the deeper layers of social meaning embedded within images. We saw in our findings that as the representational category became “thicker” (e.g., moving from “incorrectness” to “connotation”), the evaluation process became more subjective and reliant on shared cultural understanding. Yet the field of AI is increasingly using thin metrics to evaluate thick concepts—i.e., using metrics (such as accuracy) and benchmarks that are better suited for evaluations of the physical world to instead evaluate the representation of social worlds. We thus advocate for an evaluation ecosystem that creates congruence between the construct (i.e., cultural representation, or its dimensions) and the measurement method, which complements thin evaluations with thick evaluations. In the words of Jackson, “[T]hin description is the necessary starting point for social investigation but not nearly enough all by itself” (Jackson, 2013), or, as Ryle argues, “thick description is a many layered sandwich, of which only the bottom slice is catered for by the thinnest description” (Ryle, 1968).
5.2. Co-Constructing Evaluations of AI’s Cultural Representation with Communities
@@
Our study shows that there is no objective conceptualization of representation to optimize for in images. What should ideally appear when a model is prompted for “Indian,” “Pakistani,” or “American” is a culturally- and contextually-contingent choice. Thus, most representational evaluations require evaluators to make interpretations, involving culturally-situated judgment, rooted in social experience. Who gets to decide what representation is and how a culture should be depicted? Who has the power to shape meaning in the struggle over meaning-making that is representation in media (cf. Desai, 2000; Hall, 1997; Abu-Lughod, 1991)? Currently, the power of this choice lies with AI researchers and anonymous evaluators who often have limited experience with the social worlds subject to evaluation. We argue, as with AI fairness work, that we must “bring the people back in” to AI measurement of representation (Denton et al., 2020; D’ignazio and Klein, 2023), to make it more contextual and remove translation gaps between abstract metrics and the actual risks and harms they seek to measure (Jacobs and Wallach, 2021).
@@
Fairness literature has argued for the importance of including stakeholder perspectives in the design of fairness metrics (Smith et al., 2024). While prior work in AI has focused on qualitative evaluations of representation, they have primarily sought human participation in evaluations to compare differences of judgment across groups and against gold standards or to distill people’s own definitions of representation. These approaches center on stakeholders’ judgments with less focus on differences in the underlying evaluative processes that inform those judgments. As an example, disagreement itself is a robust area of research in NLP and AI, with much focus on the sociocultural influences that shape human judgment (Davani et al., 2022; Díaz et al., 2022). Studying the culturally-situated ways that people evaluate and understand representation in our study allowed us to suggest metrics and measures of representation that are more in line with the ways people who actually experience and consume images will measure their own representation. Thus, our findings echo researchers like Smith et al. (2024) who call for AI measurement to come in line with the actual experiences of people on the ground—not just the experiences of researchers—and integrate their experiences, knowledge, and expertise in the measurement process. The aim is not to simply apply the same metrics but to co-construct metrics and measures with them. Integration of expertise for representation must go beyond large-scale, globally-deployed surveys often used by AI researchers. This approach offers only a very limited mechanism to integrate cultural knowledge from particular communities into evaluative frameworks, because the metrics themselves are developed by AI researchers. As prior work argues, annotation evaluation pipelines are structured to align resulting data with what data requesters define to be ground truth (Miceli et al., 2020; Posada, 2023; Chandhiramowuli et al., 2024). As a result, evaluators must constrain their assessments within the feedback categories set by researchers. Evaluations outside of these constraints for evaluation tasks are conflated with other concerns, or dismissed altogether.
@@
Building on work on subjectivity in annotation (Cabitza et al., 2023; Davani et al., 2022) and work in HCI that contests the idea of singular ground truth (Muller et al., 2021; Gordon et al., 2022), our study suggests opportunities for developing evaluation methods for representation that leverage discursivity and deliberation (cf. Shen et al., 2022). Bergman et al. (2024) have highlighted the value of discursive deliberation in eliciting individuals’ justifications for the views they espouse. To do this, evaluative methods might actively encourage collective conversation and disagreement among evaluators through interpretive methods like workshops in addition to thin methods like benchmarks. In our study, discursive evaluation helped to fill knowledge gaps, such as whether people in different regions celebrated certain occasions in the same way. However, our discursive evaluations also revealed and helped to resolve disagreements that were ontological in nature, about the very nature of representation. Ontological disagreement here was not simply a matter of knowledge gaps, but rather fundamental differences in understanding what a depiction represents and what it should represent. Discursive evaluation processes both shed light on ontological disagreements and shift understandings in real time. Thus, dialogue and discussion can change or reaffirm evaluative judgments evaluators might have held before—changing the nature of feedback evaluators might give in individual annotation tasks.
5.3. Making Space for Thickness in AI Practice
@@
As we build the case for thick evaluations for cultural representations, we also reflect on the epistemological hurdles that will emerge if thick evaluations were to be adopted by AI developers. For one, thick evaluations are at odds with values like generalization and scale in computing (Birhane et al., 2022; Hanna et al., 2020). Thick evaluations demand a deep engagement with specific cultural contexts, requiring time and investment, and acknowledging the diversity of interpretations and meanings associated with representation. They embrace subjectivity and multiple interpretations, acknowledging that there is no single “correct” way to represent a social world. This all limits their scalablity and generalizability—as a result, they also empirically demonstrate the limits of scale and generalizability. As Tsing reminds us, the multiplicity of social worlds resists attempts at hegemonic structuring (Tsing, 2012).
@@
For representational evaluations to become ‘thicker,’ AI practitioners and researchers will also have to stop drawing on the epistemic cultures of positivism and techno-solutionism prevalent in ML and computing (Morozov, 2013; Lindtner et al., 2016; Cunningham et al., 2023). These epistemologies in AI have been extensively critiqued in AI fairness. For instance, annotation practices rarely account for social histories of race and gender (Scheuerman et al., 2020), and more broadly, algorithms fail to capture social nuances and lived realities (e.g., Broussard, 2018; Mattern, 2021). The critiques of thin-ness our findings proffer are also in line with critiques of positivism in the social sciences. For instance, Babones argues against treating humans and society as a “knowable objective reality, represented reasonably well by the variables” (Babones, 2016) and Kitchin argues this work produces social analysis that “is reductionist, functionalist and ignores the effects of culture, politics, policy, governance and capital” (Kitchin, 2014). In this paper, we demonstrate how current approaches to evaluating representation thus ignore extensive critiques of quantitative measurement in the social sciences, which caution against attempting to formalize fundamentally social processes, like representation, into quantifiable metrics or objective functions that assume one base empirical reality (Steinmetz, 2005). Our findings demonstrate empirically how limited the prevailing paradigms of objectivity and quantifiability are when faced with evaluating social worlds, and how much richer AI evaluations could be if we embraced the pluralism of thick evaluations, with their emphasis on situated context, subjectivity, and lived experiences. Such a move would require de-centering objectivity, acknowledging the situatedness of evaluation and cultural representation.
@@
Our findings also show the need for a fundamental shift in how we understand and value knowledge in AI practice. The practice of thickness needs communities’ participation in co-constructing evaluations. That is why in this paper we draw on interpretive, dialogue-based methods to capture the discursive processes involved in negotiations of the meaning of images. Our methods also recognized the expertise brought by local experts and stakeholders who could partner with us to create a sample situated in the diverse social worlds of participants. However, adopting this practice would require AI researchers to recognize and embrace pluralistic ways of knowing and understanding the world. Integrating qualitative, situated knowledge into the predominantly quantitative realm of AI necessitates a shift in mental models and a confrontation with deeply ingrained assumptions about considerations of expertise (cf. Madaio et al., 2024; Díaz and Smith, 2024).
We take inspiration from calls like those of Agre who called for a critical technical practice around computing (Agre, 2014), to call for a practice of thickness not just in measurement but also in computing writ large. This would require questioning the singular focus on scale and generalization, embracing qualitative methods and epistemologies, actively co-creating AI with communities and reflexively interrogating the epistemological values embedded in our evaluation practices.
6. Conclusion
@@
As AI technologies are being proposed to engage in cultural production, and research continues to highlight the failures of AI to adequately represent all cultures, our work suggests the need to develop new, thicker forms of evaluations of cultural representation, to better reflect how people consume and interpret images of their cultures. Thus, in this paper, we introduce a framework for thick evaluations of cultural representation, developed in conversation with participants in South Asia, that can move the field beyond existing reductive ideals of representation. Showing that representation is a process of meaning-making through the interaction between image, viewer and context, we argue that measurement of representation in AI needs to encompasses more than just factual correctness, creating practices that bridge the disconnect between constructing representation as a technical goal and understanding it as a social goal. Developing congruence between the thickness of representation as a concept and thickness of the evaluation method will ensure AI practitioners avoid creating brittle, impoverished evaluations and mitigations for representation in the design of AI systems, and, ideally, lead towards AI systems that are better able to represent the plurality of social worlds of diverse cultures.

### v2

@@ Discussion
In this discussion section, we reflect on three implications of our findings for the practice of AI evaluations: (1) the need to reconsider the construct of representation underpinning AI evaluations by creating congruence between that construct and its measurement, (2) the importance of co-constructing methods for evaluating representation with communities, and (3) the tensions inherent in attempts to “thicken” evaluation practices for cultural representation in AI.
Thick Methods for Thick Constructs of Cultural Representation
@@
Our findings highlight the urgent need for a critical re-appraisal of how we conceptualize and operationalize representation in AI evaluations. Instead of representation being a singular fixed concept or objective truth, participants interpreted it in situated and negotiated ways via multiple granular categories. Our participants empirically demonstrated that what gets bundled into the ‘suitcase word’ of representation (cf. Chasalow and Levy 2021) are multiple unobservable constructs of cultural representation, each requiring different approaches to measurement and metrics. Our findings thus suggest that AI evaluations typically conflate two constructs of representation—one that is thinner, which focuses on empirical accuracy of physical worlds, and one that is thicker, such as “coherence,” which reflects the internal social consistency and plausibility of the social world depicted in an image.
@@
While thin evaluations may be suitable for assessing “thinner” constructs like the factual incorrectness of images’ correspondence with an empirical reality or ground truth (e.g., landmarks), thicker constructs, such as those related to cultural norms, social relations, and power dynamics, necessitate thick evaluations that can capture the deeper layers of social meaning within images. We saw in our findings that as the representational category became “thicker” (e.g., moving from “incorrectness” to “connotation”), the evaluation process became more subjective and reliant on shared cultural understanding. Yet the field of AI is increasingly using thin metrics to evaluate thick concepts—i.e., using metrics and benchmarks that are better suited for evaluations of the physical world to instead evaluate the representation of social worlds (cf. Nelson 2023a, b). We thus advocate for an evaluation ecosystem that creates congruence between the construct (i.e., cultural representation, or its dimensions) and the measurement method, which complements thin evaluations with thick evaluations. In the words of Jackson (2013), “[T]hin description is the necessary starting point for social investigation but not nearly enough all by itself,” or, as Ryle (1968) argues, “thick description is a many layered sandwich, of which only the bottom slice is catered for by the thinnest description.”
Co-Constructing Evaluations of AI’s Cultural Representation with Communities
@@
Our study shows that there is no objective conceptualization of representation to optimize for in images. What should ideally appear when a model is prompted for “Indian,” “Pakistani,” or “American” is a culturally- and contextually-contingent choice. Thus, most representational evaluations require evaluators to make interpretations, involving culturally-situated judgment, rooted in social experience. Who gets to decide what representation is and how a culture should be depicted? Who has the power to shape meaning in the struggle over meaning-making that is representation in media (cf. Desai 2000; Hall 1997; Abu-Lughod 1991)? Currently, the power of this choice lies with AI researchers and anonymous evaluators who often have limited experience with the social worlds subject to evaluation. We argue, as with AI fairness more broadly, that we must “bring the people back in” to AI measurement of representation (Denton et al. 2020; Smith et al. 2024), to make it more contextual and remove translation gaps between abstract metrics and the actual risks and harms they seek to measure (Jacobs and Wallach 2021; D’ignazio and Klein 2023).
@@
Although some prior work has conducted qualitative evaluations of representation in AI, they have primarily focused on on differences in stakeholders’ judgments or decisions compared to some gold standard, with less focus on differences in the underlying evaluative processes that inform those judgments. As an example, disagreement itself is a robust area of research in NLP and AI, with much focus on the sociocultural influences that shape human judgment (Davani et al. 2022; Díaz et al. 2022). Studying the culturally-situated ways that people evaluate and understand representation in our study allowed us to suggest metrics and measures of representation that are more in line with the ways people who actually experience and consume images interpret representation of their culture. Thus, our findings echo researchers like Smith et al. (2024) who call for AI measurement to come in line with the actual experiences of people on the ground—not just the experiences of researchers—and integrate their experiences, knowledge, and expertise (cf. Díaz and Smith 2024) in the measurement process. The aim is not simply to apply the same metrics in different ways, but to co-construct metrics and measures with stakeholders. Integration of expertise for representation must go beyond large-scale, globally-deployed surveys often used by AI researchers. This approach offers only a very limited mechanism to integrate cultural knowledge from particular communities into evaluative frameworks, because the metrics themselves are developed by AI researchers. Annotation evaluation pipelines are structured to align resulting data with what data requesters define to be ground truth (Miceli et al. 2020; Posada 2023; Chandhiramowuli et al. 2024; Muller et al. 2021). As a result, annotators or evaluators must constrain their assessments within the feedback categories set by researchers. Evaluations outside of these constraints for evaluation tasks are conflated with other concerns, or dismissed altogether (Posada 2023).
@@
Building on work on subjectivity in annotation (Cabitza et al. 2023; Davani et al. 2022) and work in HCI that contests the idea of singular ground truth (Muller et al. 2021; Gordon et al. 2022), our study suggests opportunities for developing evaluation methods for representation that leverage discursiveness and deliberation (cf. Shen et al. 2022). Bergman et al. (2024) have highlighted the value of discursive deliberation in eliciting individuals’ justifications for the views they espouse. To do this, evaluative methods might actively encourage collective conversation and disagreement among evaluators through interpretive methods like workshops in addition to thin methods like benchmarks. In our study, discursive evaluation helped to fill knowledge gaps, such as whether people in different regions celebrated certain occasions in the same way. However, our discursive evaluations also revealed and helped to resolve disagreements that were ontological in nature, about the very nature of representation. Ontological disagreement here was not simply a matter of knowledge gaps, but instead fundamental differences in understanding what an image represents and what it should represent. Discursive evaluation processes both shed light on ontological disagreements and shift understandings in real time. Thus, dialogue and discussion can change or reaffirm evaluative judgments evaluators might have held before—changing the nature of feedback evaluators might give in individual annotation tasks.
Making Space for Thickness in AI Practice
@@
As we build the case for thick evaluations for cultural representations, we also reflect on the epistemological hurdles that may emerge if thick evaluations were to be adopted by AI developers. For one, thick evaluations are at odds with incentives and values like generalization and scale in computing (Birhane et al. 2022b; Hanna et al. 2020). Thick evaluations demand a deep engagement with situated cultural contexts (cf. Arzberger et al. 2024), requiring time and investment, and acknowledging the diversity of interpretations and meanings associated with representation. They embrace subjectivity and multiple interpretations, acknowledging that there is no single “correct” way to represent a social world. This all limits their scalablity and generalizability—but at the same time, they also empirically demonstrate the limits of scale and generalizability. As Tsing (2012) reminds us, the multiplicity of social worlds resists attempts at hegemonic structuring into scalable units.
@@
For representational evaluations to become ‘thicker,’ AI practitioners and researchers will also have to stop drawing on the epistemic cultures of positivism and techno-solutionism prevalent in ML and computing (Morozov 2013; Lindtner et al. 2016; Cunningham et al. 2023). These epistemologies in AI have been extensively critiqued in AI fairness. For instance, annotation practices rarely account for social histories of race and gender (Scheuerman et al. 2020), and more broadly, algorithms fail to capture social nuances and lived realities (e.g., Broussard 2018; Mattern 2021; Arzberger et al. 2024). The critiques of thin-ness our findings proffer are also in line with critiques of positivism in the social sciences. For instance, Babones (2016) argues against treating humans and society as a “knowable objective reality, represented reasonably well by the variables” and Kitchin (2014) argues that positivism produces social analysis that “is reductionist, functionalist and ignores the effects of culture, politics, policy, governance and capital.” In this paper, we demonstrate how current approaches to evaluating representation thus ignore extensive critiques of quantitative measurement in the social sciences, which caution against attempting to formalize fundamentally social processes, like cultural representation, into quantifiable metrics or objective functions that assume a singular empirical reality (Steinmetz 2005). Our findings demonstrate how limited the prevailing paradigms of objectivity and quantifiability are when attempting to evaluate social worlds, and how much richer AI evaluations could be if we embraced the pluralism of thick evaluations, with their emphasis on situated context, subjectivity, and lived experiences. Such a move would require de-centering objectivity, acknowledging the situatedness of evaluations of cultural representation.
@@
Our findings also show the need for a fundamental shift in how we understand and value knowledge in AI practice. The practice of thickness needs communities’ participation in co-constructing evaluations. That is why in this paper we draw on interpretive, dialogue-based methods to capture the discursive processes involved in negotiations of the meaning of images. Our methods also recognized the expertise brought by local experts and stakeholders who could partner with us to create a sample situated in the diverse social worlds of participants. However, adopting this practice would require AI researchers to recognize and embrace pluralistic ways of knowing and understanding the world. Integrating qualitative, situated knowledge into the predominantly quantitative realm of AI necessitates a shift in mental models and a confrontation with deeply ingrained assumptions about considerations of expertise (cf. Madaio et al. 2024; Díaz and Smith 2024). We take inspiration from those like Agre who call for a critical technical practice around computing (Agre 2014; Malik and Malik 2022), and those like Nelson (2023a, b) to call for a practice of thickness not just in measurement but also in AI writ large. This would require questioning the singular focus on scale and generalization, embracing qualitative methods and epistemologies, actively co-creating AI with communities and reflexively interrogating the epistemological values embedded in our evaluation practices.
Limitations and Future Work
This work has several limitations, which future work may address. First, our participants were from three countries in South Asia; bringing their own cultural and individual perspectives on the most salient elements for which to evaluate representation. Since generalizability is not the goal of interpretive qualitative research, which instead aims for transferability, (Soden et al. 2024; Drisko 2025), we do not make any claims of generalizability of this research to other locales. Thus, future research could help shed light on the extent to which similar taxonomies of thick evaluations may be transferrable (Drisko 2025) or extendable for representational evaluations with different contexts or populations. In addition, we focused on text-to-image AI systems in this study. Although we believe our work could speak to thick evaluations of representation in other AI modalities (e.g., text, video, audio, or multi-modal systems), it is an open question of precisely how such thick evaluations in other modalities ought to be conducted, or how the approach ought to change, if so. In this work, we discuss how certain elements of cultural representation may be best evaluated by thick evaluations, while others (e.g., the “incorrectness” dimension) may be better suited to thin evaluations, or correspondence to a ground truth; however, future work should be conducted to validate this hypothesis, to better understand which elements of cultural representation are best evaluated with which approaches. Finally, although in this paper, we map out a space for conducting thick evaluations, future work should explore how best to incorporate the results of these evaluations into shaping the design and development of generative AI systems, in order to avoid the failures of cultural representation that motivated this work.
Conclusion
@@
As AI technologies are being proposed to engage in cultural production, and research continues to highlight the failures of AI to adequately represent all cultures, our work suggests the need to develop new, thicker forms of evaluations of cultural representation, to better reflect how people consume and interpret images of their cultures. Thus, in this paper, we introduce a framework for thick evaluations of cultural representation, developed in conversation with participants in South Asia, that can move the field beyond existing reductive ideals of representation. Showing that representation is a process of meaning-making through the interaction between image, viewer, and context, we argue that measurement of representation in AI needs to encompasses more than just factual correctness, creating practices that bridge the disconnect between constructing representation as a technical goal and understanding it as a social goal. Developing congruence between the thickness of representation as a concept and thickness of the evaluation method will ensure AI practitioners avoid creating brittle, impoverished evaluations and mitigations for representation in the design of AI systems, and, ideally, lead towards AI systems that are better able to represent the plurality of social worlds of all cultures.
Appendix
We include here a table of detailed demographic information of our participants from our study. We present their details by countries, please see the tables below:
Participants Country Age Gender Educational Level Region Language
LK-1 Sri Lanka 25–34 Male Undergraduate North Western province Sinhala
LK-2 Sri Lanka 18–24 Male Undergraduate North Western Province Sinhala
LK-3 Sri Lanka 18–24 Female Undergraduate Western Province Tamil, English
LK-4 Sri Lanka 18–24 Male Secondary Western Province Sinhala
LK-5 Sri Lanka 18–24 Female Undergraduate Western Province Sinhala
LK-6 Sri Lanka 18–24 Female Undergraduate Western Province/ Central Province Sinhala
LK-7 Sri Lanka 18–24 Female Undergraduate Central Province/ Western Province Sinhala,Tamil and English
LK-8 Sri Lanka 25–34 Male Undergraduate Western Province/ North Western Province Tamil
LK-9 Sri Lanka 18–24 Male Undergraduate Western Province/ Southern Province/ Central province Sinhala
LK-10 Sri Lanka 35–44 Male Undergraduate Western Province Sinhala
LK-11 Sri Lanka 45–54 Female Secondary Western Province Sinhala
Table 1: Participant Information from Sri Lanka
Participants Country Age Gender Educational Level Region Language
PK-1 Pakistan 25–34 Female Post-Graduate Punjab Urdu
PK-2 Pakistan 25–34 Male Undergraduate Gilgit-Baltistan Balti
PK-3 Pakistan 25–34 Male Post-Graduate Rajanpur Balochi, Siraiki, Urdu
PK-4 Pakistan 35–44 Transgender Man Undergraduate Punjab Urdu
PK-5 Pakistan 18–24 Transgender woman Undergraduate Punjab Punjabi, Urdu
PK-6 Pakistan 18–24 Female Undergraduate Gilgit Burushaski
PK-7 Pakistan 18–24 Female Secondary Punjab Urdu, Punjabi
PK-8 Pakistan 18–24 Female Undergraduate Punjab Urdu, Punjabi
PK-9 Pakistan 18–24 Male Secondary Gilgit Baltistan Balti
PK-10 Pakistan 18–24 Male Secondary Gilgit Baltistan Balti
PK-11 Pakistan 18–24 Male Undergraduate South Punjab Saraiki
Table 2: Participant Information from Pakistan
Participants Country Age Gender Educational Level Region Language
IN-1 India 25–34 Female Post-Graduate Kerala Malayalam
IN-2 India 55–64 Female Undergraduate Maharastra/ Karnataka English, French, Hindi
IN-3 India 25–34 Female Undergraduate Assam Assamese and Hindi
IN-4 India 45–54 Female Post-Graduate Maharashtra Marathi
IN-5 India 45–54 Female Post-Graduate Chandigarh/ Punjab Punjabi, Hindi, English
IN-6 India 25–34 Male Post-Graduate Bihar Hindi
IN-7 India 18–24 Male Secondary Gujarat Gujarati
IN-8 India 25–34 Non-binary Post-Graduate Maharashtra Hindi, Magadhi
IN-9 India 25–34 Genderqueer Post-Graduate West Bengal, Gujarat, Karnataka Bengali and English
IN-10 India 25–34 Female Undergraduate Telegana/ Karnataka Telugu, Hindi, Kannada
IN-11 India 18–24 Male Undergraduate Rajasthan Rajasthani
IN-12 India 18–24 Male Undergraduate Chhattisgarh Chhattisgarhi [^1] Chhattisgarh local Language mix with Hindi known as Chhattisgarhi
IN-13 India 25–34 Male Post-Graduate Tripura Sylheti
IN-14 India 25–34 Female Post-Graduate Assam Assamese
IN-15 India 35–44 Female Post-Graduate Kashmir Kashmiri, Urdu, Hindi, English
Table 3: Participant Information from India